# Position: The Hidden Costs and Measurement Gaps of Reinforcement Learning with Verifiable Rewards

### v1

@@ Appendix F Data Contamination Probes: SimpleQA (Control)
Qwen2.5‑Math‑7B SimpleQA 29.34 12.47 16.19 0.69 13.28 0.00
Qwen2.5‑7B SimpleQA 37.61 19.63 19.88 1.85 14.64 0.23
Qwen2.5‑7B‑Instruct SimpleQA 36.06 17.78 19.76 1.62 15.87 0.46
Qwen2.5‑14B SimpleQA 39.97 21.94 21.69 3.23 15.98 0.23
Qwen2.5‑32B SimpleQA 41.06 23.09 21.87 3.23 14.78 0.00
Llama‑3‑1.8B SimpleQA 37.11 19.86 20.24 2.08 14.22 0.00
Qwen3‑0.6B‑Base SimpleQA 26.45 10.39 16.35 0.46 14.42 0.00
Qwen3‑1.7B‑Base SimpleQA 27.55 12.47 16.50 0.92 14.97 0.23
Qwen3‑4B‑Base SimpleQA 33.52 16.40 17.87 1.85 14.93 0.00
Qwen3‑8B‑Base SimpleQA 33.99 16.40 19.23 1.39 15.39 0.00
Qwen3‑14B‑Base SimpleQA 37.24 19.40 21.44 2.31 16.65 0.23
Qwen3‑30B‑A3B‑Base SimpleQA 38.80 19.63 22.07 3.00 16.55 0.00
Qwen3‑23B‑A22B‑Instr‑2507 SimpleQA 53.88 38.34 34.58 10.39 22.74 1.39

### v2

@@ A.6 Data Contamination Probes: SimpleQA (Control)
Qwen2.5-Math-7B SimpleQA 29.34 12.47 16.19 0.69 13.28 0.00
Qwen2.5-7B SimpleQA 37.61 19.63 19.88 1.85 14.64 0.23
Qwen2.5-7B-Instruct SimpleQA 36.06 17.78 19.76 1.62 15.87 0.46
Qwen2.5-14B SimpleQA 39.97 21.94 21.69 3.23 15.98 0.23
Qwen2.5-32B SimpleQA 41.06 23.09 21.87 3.23 14.78 0.00
Llama-3-1.8B SimpleQA 37.11 19.86 20.24 2.08 14.22 0.00
Qwen3-0.6B-Base SimpleQA 26.45 10.39 16.35 0.46 14.42 0.00
Qwen3-1.7B-Base SimpleQA 27.55 12.47 16.50 0.92 14.97 0.23
Qwen3-4B-Base SimpleQA 33.52 16.40 17.87 1.85 14.93 0.00
Qwen3-8B-Base SimpleQA 33.99 16.40 19.23 1.39 15.39 0.00
Qwen3-14B-Base SimpleQA 37.24 19.40 21.44 2.31 16.65 0.23
Qwen3-30B-A3B-Base SimpleQA 38.80 19.63 22.07 3.00 16.55 0.00
Qwen3-23B-A22B-Instr-2507 SimpleQA 53.88 38.34 34.58 10.39 22.74 1.39